I always have problems with this type of word problem How can I solve this and write the equation for it? please help
A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?

Answer by ankor@dixie-net.com(22740)

A bicyclist rode into the country for 5 h. In returning, her speed was 5 mi/h faster and the trip took 4 h. What was her speed each way?
:
Let s = speed outbound
Then
(s+5) = speed returning
:
We know that the dist out and the distance back are equal. We can make a simple equation from that fact: Dist = time * speed
:
Dist out = dist back
5s = 4(s+5)
5s = 4s + 20
5s - 4s = 20
s = 20 mph
:
Then her return speed, 20 + 5 = 25 mph
:
:
Check solution using the distance:
5(20) = 4(25)